A radiator contains 12 quarts of fluid, 15% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 35% antifreeze? (Round your answer to two decimal places.)

Answer by lwsshak3(11628)

A radiator contains 12 quarts of fluid, 15% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 35% antifreeze? (Round your answer to two decimal places.)
**
let x=amt of fluid to be drained and replaced with pure antifreeze.
12-x=amt of fluid remaining in radiator
..
15%(12-x)+100%x=35%*12
1.8-.15x+x=4.2
.85x=2.4
x=2.4/.85
x=2.82
amt of fluid to be drained and replaced with pure antifreeze=2.82 quarts